Displacement 20 40 40 80 80 60 40 0
Time 10 20 30 40 50 60 70 80
Explain the different sections of the graphs as much as details as you can
1
Expert's answer
2019-12-11T10:22:08-0500


From 10 s. to 20 s.: motion frontwards with velocity:


"\frac{40-20}{10}=2\frac{m}{s}"

From 20 s. to 30 s.: no motion.

From 30 s. to 40 s.: motion frontwards with velocity:


"\frac{80-40}{10}=4\frac{m}{s}"

From 40 s. to 50 s.: no motion.

From 50 s. to 60 s.: motion backwards with velocity:


"\frac{60-80}{10}=-2\frac{m}{s}"

From 60 s. to 70 s.: motion backwards with velocity:


"\frac{40-60}{10}=-2\frac{m}{s}"

From 70 s. to 80 s.: motion backwards with velocity:


"\frac{0-40}{10}=-4\frac{m}{s}"